Possible outcomes for three investment alternatives and their probabilities of occurrence are given next.       
  

Alternative 1 Alternative 2 Alternative 3
Outcomes Probability Outcomes Probability Outcomes Probability
  Failure 50 .2 90 .3 95 .2
  Acceptable 90 .4 190 .3 215 .6
  Successful 135 .4 225 .4 380 .2

  
Using the coefficient of variation, rank the three alternatives in terms of risk from lowest to highest. (Do not round intermediate calculations. Round your answers to 3 decimal places.)
Answer #1
Alternative-1
Mean return expected=50*0.2+90*0.4+135*0.4 100.00
Variance=0.2*(50-100)^2+0.4*(90-100)^2+0.4*(135-100)^2 1030.00
Standard Deviation=Square Root of Variance=(1030)^(1/2) 32.09
coefficient of variation=32.09/100 0.32
Alternative-2
Mean return expected=90*0.3+0.3*190+0.4*225 174.00
Variance=0.3*(90-174)^2+0.3*(190-174)^2+0.4*(225-174)^2 3234.00
Standard Deviation=Square Root of Variance=(3234)^(1/2) 56.87
Coefficient of variation=56.87/174 0.33
Alternative-3
Mean return expected=0.2*95+0.6*215+0.2*380 224.00
Variance=0.2*(95-224)^2+0.6*(215-224)^2+0.2*(380-224)^2 8244.00
Standard Deviation=Square Root of Variance=(8244)^(1/2) 91
Coefficient of variation=91/224 0.41
Ranks are as follows:
Alternative-1 : rank 1
Alternative-2 : rank 2
Alternative-3 : rank 3